# Rethinking Graph Structure Learning in the Era of LLMs

### v3

@@ Abstract
Recently, the emergence of LLMs has prompted researchers to integrate language descriptions into graphs, aiming to enhance model encoding capabilities from a data-centric perspective. This graph representation is called text-attributed graphs (TAGs). A review of prior advancements highlights that graph structure learning (GSL) is a pivotal technique for improving data utility, making it highly relevant to efficient TAG learning. However, most GSL methods are tailored for traditional graphs without textual information, underscoring the necessity of developing a new GSL paradigm. Despite clear motivations, it remains challenging: (1) How can we define a reasonable optimization objective for GSL in the era of LLMs, considering the massive parameters in LLM? (2) How can we design an efficient model architecture that enables seamless integration of LLM for this optimization objective? For Question 1, we reformulate existing GSL optimization objectives as a tree optimization framework, shifting the focus from obtaining a well-trained edge predictor to a language-aware tree sampler. For Question 2, we propose decoupled and training-free model design principles for LLM integration, shifting the focus from computation-intensive fine-tuning to more efficient inference. Based on this, we propose Large Language and Tree Assistant (LLaTA), which leverages tree-based LLM in-context learning to enhance the understanding of topology and text, enabling reliable inference and generating improved graph structure. Extensive experiments on 10 datasets demonstrate that LLaTA enjoys flexibility-incorporated with any backbone; scalability-outperforms other LLM-enhanced graph learning methods; effectiveness-achieves SOTA predictive performance.
@@ 1 Introduction
To achieve this, graph structure learning (GSL) is a promising approach that enhances data utility for vanilla graph neural networks (GNNs). However, most existing prevalent GSL approaches [35] are designed for traditional graphs and fail to process rich textual information, leading to sub-optimal performance. Despite significant efforts by recent LLM-based GSL methods, such as GraphEdit [7] and LLM4RGNN [31], their optimization objectives and model architectures still adhere to traditional GSL paradigms, inheriting unnecessary complexities. Specifically, these prominent GSL approaches [32, 28, 10, 7, 31] rely on carefully designed loss functions to dynamically maintain a graph learner (i.e., structure optimizer) coupled with well-defined instruction datasets or a specific GNN backbone (i.e., improved structure tailored for the specific downstream encoder). Based on this, we have the following 2 observations and seek to explore them to gain key insights into developing a new GSL paradigm in the era of LLMs.
@@
Answer 1: In Sec. 3.1, we present a new optimization framework that reformulates the existing GSL optimization objective (well-trained edge predictor) as a tree-based optimization task (well-defined language-aware tree sampler).
@@
Based on the aforementioned key insights from Answers, we propose Large Language and Tree Assistant (LLaTA) as follows: (1) Topology-aware In-context Construction: We first quantify the dynamic complexity of graph topology via structural entropy. By employing the greedy algorithm to minimize this measurement, we construct a hierarchical structural encoding tree, which captures topology insights centered on multi-level communities (i.e., non-leaf nodes within the tree). (2) Tree-prompted LLM Inference: Subsequently, this tree serves as a high-quality prompt, enabling LLMs to perform in-context learning for a comprehensive understanding of topology and text. Specifically, we utilize LLM to reveal textual semantic relationships within the leaf community. Based on these topology and text insights, we reallocate the leaf dependency to achieve tree optimization. (3) Leaf-oriented Two-step Sampling: Finally, we perform an LLM-guided leaf selection to identify the current node and candidate nodes for edge addition or removal, enabling training-free GSL. The core idea of LLaTA is to enable LLM in-context learning through topology-aware tree prompts, eliminating the need for expensive fine-tuning. In Sec. 3.3, we provide the empirical investigation, demonstrating the effectiveness of this new GSL paradigm in the era of LLMs.
Our contributions. (1) New Perspective. We systematically review existing GSL methods and provide empirical investigations, revealing the challenges and opportunities for GSL in the era of LLMs. (2) Innovative Approach. We reformulate the existing GSL and propose a tree-based optimization framework with model design principles. Based on this, we introduce LLaTA, which seamlessly integrates both topology and text insights by tree-driven prompts to facilitate LLM in-context learning and generate improved graph structure, with complete theoretical support. (3) SOTA Performance. Extensive experiments demonstrate the superior performance of LLaTA. It outperforms recent LLM-based methods by 1.3%-2.5% in accuracy while running 3.0h-9.7h faster.
@@ 2.1 Notations and Problem Formulation
Graph Structure Learning in TAGs. In this paper, we aim to improve the graph structure $\mathbf{A}^{\star}$ for any downstream task by incorporating original topology and textual information. Given the output predictions $\hat{\mathbf{Y}}$ from the downstream backbone with improved structure, the general loss is formulated as $\mathcal{L}_{\text{task }}(\hat{\mathbf{Y}},\mathbf{Y})+\alpha\mathcal{L}_{\text{reg }}\left(\mathbf{A}^{\star},\mathbf{A}\right)$, where $\mathcal{L}_{\text{task }}$ evaluates specific task performance (e.g., node classification), and $\mathcal{L}_{\text{reg }}$ establishes the guiding principles for improved structure.
@@ 2.3 Related Works
SE-based Methods. Existing SE-based methods leverage the well-defined structural encoding tree to perform a wide range of graph-based downstream tasks such as node clustering [22], community detection [17], multi-layer coarsening [27], and embedding dimension estimation [30]. Recent SEGSL [37] enhances edge connectivity and quality using this structural encoding tree, but it relies on end-to-end tree-based training without incorporating LLMs.
@@
LLM-based GSL. GraphEdit [7] improves node connectivity by adding neighbors through a well-trained edge predictor and refines the improved graph structure by evaluating node relevance using a fine-tuned LLM. LLM4RGNN [31] leverages a fine-tuned LLM to infer node relevance, which is used to train an edge predictor. Compared to improving the graph structure, it focuses more on robustness against graph adversarial attacks.
@@ 3.1 GSL Optimization Objectives Reformulation
As highlighted by Obs.1 in Sec. 1, existing GSL methods aim to train an effective edge predictor as the graph learner. This typically requires coupling the graph learner with a specific backbone and jointly optimizing them. Although effective, this objective becomes inefficient for LLMs due to their large number of parameters. Our investigation shows that current LLM-based GSL approaches often adopt instruction fine-tuning to mitigate full-parameter training. However, understanding complex graph structures and crafting appropriate instruction datasets can be labor-intensive, as reflected in their elaborate workflows [7, 31]. To address this, we reformulate GSL and introduce a tree-based optimization framework that avoids edge predictors, as illustrated in Fig. 1 (a)-(c). The core idea is to develop a language-aware tree sampler for efficient GSL.
@@
Specifically, (a) Tree Construction first reorganizes the original graph to assist LLMs in understanding the topology. Based on this, (b) Tree Optimization leverages textual information and LLM to perform tree optimization. Finally, (c) Improved Structure is generated by a language-aware tree sampler. Please refer to Sec.3.2 and Sec.4 for implementation of the above framework in LLaTA.
@@ 3.2 Model Architecture Review
As outlined by Obs. 2 in Sec. 1, integrating LLMs in GSL remains a challenge. To establish model design principles for tree optimization, we revisit existing GSL paradigms.
The Coupled Paradigm tightly integrates the graph learner and backbone for task-specific learning. However, it faces limitations such as unstable performance due to restricted generalizability and fairness concerns, and inflexible deployment as switching tasks or backbones requires retraining. In contrast, the Decoupled Paradigm trains the graph learner and backbone independently, offering better compatibility with LLMs for GSL. Although recent LLM-based GSL methods adopt this approach, they still face challenges such as indirect integration of topology and text, reliance on instruction datasets, and the complexity of fine-tuning. For a more efficient and reliable approach, we recommend a fully decoupled paradigm, reducing fine-tuning reliance and focusing on high-quality in-context prompts. Detailed discussion is provided in the Appendix C.
To achieve this, we leverage the in-context learning ability of LLMs to jointly understand graph topology and node text for reliable inference. Specifically, (a) Tree Construction: we first build a structural encoding tree to generate topology-aware in-context prompts, enhancing the LLM’s understanding of graph structure. Then, (b) Tree Optimization: based on these prompts, we use LLMs to infer semantic relationships among nodes from their textual descriptions, and reassign leaf dependencies to optimize the tree structure. The resulting hierarchical tree integrates both structural and textual information, forming a strong basis for GSL. Finally, (c) Improved Structure: we apply a leaf-oriented two-step sampling to select the current node and a candidate set for edge modification, yielding an improved graph for any downstream scenario. For deeper insights into the leaf-oriented sampler within GSL, please refer to Appendix B.
@@ 3.3 Empirical Investigation
To demonstrate the effectiveness of our proposal in Sec. 3.1-3.2, we present empirical results in Fig. 1(d). Due to space constraints, detailed experimental settings and analysis are provided in Appendix D. In our reports, over-smoothing ($\downarrow$) reflects the distinguishability of node embeddings, while over-squashing ($\uparrow$) indicates the ability of nodes to perceive distant ones. These metrics are widely used in recent GSL studies to quantify the quality of improved structure, as they not only directly impact downstream accuracy ($\uparrow$) but also evaluate method robustness. They demonstrate that LLM-based GSL methods generate higher-quality structures than traditional ones, underscoring the necessity of LLMs. Furthermore, LLaTA achieves the best performance, validating the effectiveness of our proposed new GSL paradigm.
@@ 4.2 Topology-aware In-context Construction
To establish a hierarchical encoding tree that simulates the natural evolution of graph topology, we constrain the height of this tree to $K$ and minimize $K$-dimensional SE proposed by [13]. The optimization objective definition:
@@
where $\operatorname{vol}(\mathcal{G})$ is the sum of the degrees of all nodes in $\mathcal{G}$, $\operatorname{vol}(\phi)$ is the sum of the degrees of nodes in $\phi$, $\phi^{+}$ is the parent of $\phi$, $\lambda$ is the root node, $\mathcal{T}^{\star}$ denotes the optimized tree that minimizes its structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$. We employ a greedy algorithm to minimize the structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$ and construct a hierarchical encoding tree $\mathcal{T}^{\star}$ with a predefined height $K$. The algorithm starts by initializing a tree of height 1, where each graph node is a leaf under a common root. It then iteratively combines node pairs to reduce $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$, followed by node lifting operations to further optimize the structure. The node lifting process continues until the tree height is reduced to $K$. The full algorithm and the operations involved is detailed in Appendix E.1.
@@
Drawing on Theorems 1 and  2, we infer that low-level communities in the encoding tree provide a more faithful representation of the graph’s topology, as they simultaneously preserve local connectivity and global structural constraints. Accordingly, for each leaf node $\alpha$, we assign its associated low-level community $\mathcal{C}^{\ell}(\alpha)$ as a topological context to enhance the accuracy of downstream inference by the LLMs. The proof process of Theorems 1 and  2 is provided in Appendix F.1-F.2.
@@ 4.3 Tree-prompted LLM Inference
Reception-aware Leaf Augmentation. To construct effective in-context prompts for LLMs, the quality of node-specific text is essential. Thus, we perform tree-guided text propagation and aggregation within each low-level community $\mathcal{C}^{\ell}(i)$ to achieve leaf augmentation for $\epsilon$-guaranteed $t^{\star}$. The key idea is that nodes in low-level communities typically show high connectivity and homophily, enabling efficient text sharing. Using this property, we filter and combine relevant texts for each leaf node based on a similarity threshold $\epsilon$, helping to reduce noise. This process is formally defined as:
$t^{\star}_{\alpha}=\operatorname{Concat}\left(\left\{t_{\alpha}\}\cup\{t_{j}\mid j\in\text{Top}_{k}\left(\left\{w_{\alpha\beta}\mid\beta\in\mathcal{C}^{\ell}(\alpha)\right\}\right),w_{\alpha\beta}\geq\epsilon\right\}\right),$ (2)
$w_{\alpha\beta}=\frac{\exp\left(\operatorname{sim}(x_{\alpha},x_{\beta})\right)}{\sum\limits_{\gamma\in\mathcal{C}^{l}(\alpha)}\exp\left(\operatorname{sim}(x_{\alpha},x_{\gamma})\right)},\>\>\>\>\operatorname{sim}(x_{\alpha},x_{\beta})=\frac{\sum_{i=1}^{f\text{-dim}}x_{\alpha}^{i}\cdot x_{\beta}^{i}}{\sqrt{\sum_{i=1}^{f\text{-dim}}(x_{\alpha}^{i})^{2}\sum_{i=1}^{f\text{-dim}}(x_{\beta}^{i})^{2}}}.$ (3)
Community of Thought. Leveraging the enhanced leaf nodes, we propose a community of thought (CoT) prompt mechanism, which contains the community-enhanced descriptions for each node and its community-related neighbors for more reliable LLM inference. Based on this, we enable the LLM to fully understand both the topology and text insights of the original graph, accurately infer the probabilities of a node belonging to each label class, and generate high-quality embeddings within the LLM’s semantic encoding space by softmax. These embedding are subsequently used for fine-grained tree optimization. In a nutshell, this approach facilitates efficient LLM in-context learning through the structural encoding tree-driven CoT. The above inference and encoding process can be formally defined as:
$z_{\alpha}=\operatorname{Extract}(ANS_{\alpha}),\>\>\>\>\text{ANS}_{\alpha}=\text{LLM-}\operatorname{CoT}(\mathcal{D}_{task},\>\mathcal{D}_{topo},\>\mathcal{D}_{sema},\>t^{\star}_{\alpha}),$ (4)
[…]
@@
Leaf Dependency Allocation. Although the structural encoding tree captures the topology of the original graph, another critical factor is the text-driven node attributes. Therefore, we optimize the low-level communities by reallocating leaf dependencies, guided by soft labels from the LLM, thus enhancing GSL with text insights. To achieve this, we first initialize the leaf-community clusters based on the original structural encoding tree. Next, we apply adaptive clustering using the silhouette coefficient to refine community assignments. The main goal is to enhance homophily within low-level communities, ensuring that most nodes in each cluster share the same label. Specifically, (1) we reallocate leaf nodes with minority labels to align with the dominant class in each community, and (2) we adaptively increase the number of clusters by leveraging the silhouette coefficient, allowing the formation of new communities that better capture homophily. An illustrative example is shown in Fig.2. This process yields the optimized structural encoding tree $\mathcal{T}^{\star}$, which serves as the foundation for defining the tree sampler. The adaptive clustering algorithm is detailed in Appendix E.2.

### v4

@@ Abstract
Recently, the emergence of LLMs has prompted researchers to integrate language descriptions into graphs, aiming to enhance model encoding capabilities from a data-centric perspective. This graph representation is called text-attributed graphs (TAGs). A review of prior advancements highlights that graph structure learning (GSL) is a pivotal technique for improving data utility, making it highly relevant to efficient TAG learning. However, most GSL methods are tailored for traditional graphs without textual information, underscoring the necessity of developing a new GSL paradigm. Despite clear motivations, it remains challenging: (1) How can we define a reasonable optimization objective for GSL in the era of LLMs, considering the massive parameters in LLM? (2) How can we design an efficient model architecture that enables seamless integration of LLM for this optimization objective? For Question 1, we reformulate existing GSL optimization objectives as a tree optimization framework, shifting the focus from obtaining a well-trained edge predictor to a language-aware tree sampler. For Question 2, we propose decoupled and training-free model design principles for LLM integration, shifting the focus from computation-intensive fine-tuning to more efficient inference. Based on this, we propose Large Language and Tree Assistant (LLaTA), which leverages tree-based LLM in-context learning to enhance the understanding of topology and text, enabling reliable inference and generating improved graph structure. Extensive experiments on 11 datasets demonstrate that LLaTA enjoys flexibility-incorporated with any backbone; scalability-outperforms other LLM-enhanced graph learning methods; effectiveness-achieves SOTA predictive performance.
@@ 1 Introduction
To achieve this, graph structure learning (GSL) is a promising approach to enhancing data utility for vanilla graph neural networks (GNNs). However, most existing GSL methods [39] are designed for traditional graphs and cannot effectively process rich textual information, leading to suboptimal performance. Despite recent efforts in LLM-based GSL approaches, such as GraphEdit [7] and LLM4RGNN [34], their optimization objectives and model architectures still follow traditional GSL paradigms, inheriting unnecessary complexities. Specifically, these prominent GSL methods [36, 29, 28, 34] rely on carefully designed loss functions to maintain a graph learner (i.e., structure optimizer) coupled with well-defined instruction datasets or specific GNN backbones (i.e., improved structure tailored for the downstream encoder). Based on this, we make two key observations and aim to explore them to guide the development of a new GSL paradigm in the era of LLMs.
@@
Answer 1: In Sec. 3.1, we present a new optimization framework that reformulates the existing GSL optimization objective (well-trained edge predictor) as a tree-based optimization task (well-defined language-aware tree sampler) to achieve topology-preserving structural optimization.
@@
Based on the insights from the previous section, we propose the Large Language and Tree Assistant (LLaTA) as follows: (1) Topology-aware In-context Construction: We quantify the dynamic complexity of graph topology via structural entropy. By applying a greedy algorithm to minimize this measure, we construct a hierarchical structural encoding tree that captures topology insights focused on multi-level communities (i.e., non-leaf nodes). (2) Tree-prompted LLM Inference: The tree serves as a high-quality prompt, enabling LLMs to perform in-context learning for a deeper understanding of both topology and text. Specifically, we use LLMs to uncover textual semantic relationships within the leaf community. Based on these insights, we reallocate leaf dependencies to refine the tree structure, optimizing it within the context of the existing communities. (3) Leaf-oriented Two-step Sampling: Finally, we perform LLM-guided leaf selection to identify nodes for edge addition or removal, achieving training-free GSL. The core idea of LLaTA is to facilitate LLM in-context learning through topology-aware tree prompts, eliminating the need for costly fine-tuning. In Sec. 3.3, we provide empirical results demonstrating the effectiveness of this new GSL paradigm.
Our contributions. (1) New Perspective. We systematically review existing GSL methods and provide empirical investigations, revealing the challenges and opportunities for GSL in the era of LLMs. (2) Innovative Approach. We reformulate the existing GSL and propose a tree-based optimization framework with model design principles. Based on this, we introduce LLaTA, which seamlessly integrates both topology and text insights by tree-driven prompts to facilitate LLM in-context learning and generate improved graph structure, with complete theoretical support. (3) SOTA Performance. Extensive experiments demonstrate the superior performance of LLaTA. It outperforms recent LLM-based methods by 1.3%-2.5% in accuracy while running 2.5h-9.2h faster.
@@ 2.1 Notations and Problem Formulation
[…]
@@ 2.3 Related Works
SE-based Methods. Existing SE-based methods utilize the structural encoding tree for various graph tasks, such as node clustering [21], community detection [16], multi-layer coarsening [27], and embedding dimension estimation [31]. Recent SEGSL [42] improves edge connectivity and quality with SE-based encoding tree, but it still relies on end-to-end tree-based training without incorporating LLMs.
@@
LLM-based GSL. GraphEdit [7] improves node connectivity by adding neighbors through a edge predictor and refines the structure using a fine-tuned LLM. LLM4RGNN [34] leverages a fine-tuned LLM to infer node relevance, which is used to train an edge predictor. In contrast, LangGSL [24] integrates LLMs with graph structure learning to jointly optimize node features and graph structure, improving performance across various tasks. In Appendix C.2, we detail the advantages of LLaTA over SEGSL and LLM-based methods.
@@ 3.1 GSL Optimization Objectives Reformulation
As highlighted by Obs.1 in Sec. 1, existing GSL methods aim to train an effective edge predictor as the graph learner. This typically requires coupling the graph learner with a specific backbone and jointly optimizing them. Although effective, this objective becomes inefficient for LLMs due to their large number of parameters. Our investigation shows that current LLM-based GSL approaches often adopt instruction fine-tuning to mitigate full-parameter training. However, understanding complex graph structures and crafting appropriate instruction datasets can be labor-intensive, as reflected in their elaborate workflows [7, 34]. To address this, we reformulate GSL and introduce a tree-based optimization framework that avoids edge predictors, as illustrated in Fig. 1 (a)-(c). The core idea is to develop a language-aware tree sampler for efficient GSL. Specifically, (a) Tree Construction first reorganizes the original graph to assist LLMs in understanding the topology. Based on this, (b) Tree Optimization leverages textual information and LLM to perform tree optimization. Finally, (c) Improved Structure is generated by a language-aware tree sampler. Please refer to Sec.3.2 and Sec.4 for implementation of the above framework in LLaTA.
@@ 3.2 Model Architecture Review
As outlined by Obs. 2 in Sec. 1, integrating LLMs in GSL remains a challenge. To establish model design principles for tree optimization, we revisit existing GSL paradigms. The Coupled Paradigm tightly integrates the graph learner and backbone for task-specific learning. However, it faces limitations such as unstable performance due to restricted generalizability and module dependent, and inflexible deployment as switching tasks or backbones requires retraining. In contrast, the Decoupled Paradigm trains the graph learner and backbone independently, offering better compatibility with LLMs for GSL. Although recent LLM-based GSL methods adopt this approach, they still face challenges such as indirect integration of topology and text, reliance on instruction datasets, and the complexity of fine-tuning. For a more efficient and reliable approach, we recommend a fully decoupled paradigm, reducing fine-tuning reliance and focusing on high-quality in-context prompts. Detailed discussion about the GSL paradigms is provided in the Appendix C.1.
To achieve this, we leverage the in-context learning ability of LLMs to jointly understand graph topology and node text for reliable inference. Specifically, (a) Tree Construction: we first build a structural encoding tree to generate topology-aware in-context prompts, enhancing the LLM’s understanding of graph structure. Then, (b) Tree Optimization: based on these prompts, we use LLMs to infer semantic relationships among nodes from their textual descriptions, and reassign leaf dependencies to optimize the tree structure based on the community structure. The resulting hierarchical tree integrates both structural and textual information, forming a strong basis for GSL. Finally, (c) Improved Structure: we apply a leaf-oriented two-step sampling to select the current node and a candidate set for edge modification, yielding an improved graph for any downstream scenario. For deeper insights into the leaf-oriented sampler within GSL, please refer to Appendix B.
@@ 3.3 Empirical Investigation
To demonstrate the effectiveness of our proposal in Sec. 3.1-3.2, we present empirical results in Fig. 1 (d). Due to space constraints, detailed experimental settings and analysis are provided in Appendix D. In our reports, over-smoothing ($\downarrow$) reflects the distinguishability of node embeddings, while over-squashing ($\uparrow$) indicates the ability of nodes to perceive distant ones. These metrics are commonly used in recent GSL studies to quantify the quality of improved structure, as they not only directly impact downstream accuracy ($\uparrow$) but also evaluate method robustness. They demonstrate that LLM-based GSL methods generate higher-quality structures than traditional ones, underscoring the necessity of LLMs. Furthermore, LLaTA achieves the best performance, validating the effectiveness of our proposed new GSL paradigm.
@@ 4.1 Topology-aware In-context Construction
To establish a hierarchical encoding tree that simulates the natural evolution of graph topology, we constrain the height of this tree to $K$ and minimize $K$-dimensional SE proposed by [12, 42]. The optimization objective definition:
@@
where $\operatorname{vol}(\mathcal{G})$ denotes the total degree of $\mathcal{G}$, $g_{\phi}$ is the sum of edges crossing the tree node, $\operatorname{vol}(\phi)$ is the sum of the degrees of nodes in $\phi$, $\phi^{+}$ is the parent of $\phi$, $\lambda$ is the root node, $\mathcal{T}^{\star}$ denotes the optimized tree that minimizes its structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$. We employ a greedy algorithm to minimize the structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$ and construct a hierarchical encoding tree $\mathcal{T}^{\star}$ with a predefined height $K$. The algorithm starts by initializing a tree of height 1, where each graph node is a leaf under a common root. It then iteratively combines node pairs to reduce $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$, followed by node lifting operations to further optimize the structure. The node lifting process continues until the tree height is reduced to $K$. The full algorithm and the operations involved is detailed in Appendix E.1.
@@
Drawing on Theorems 1 and  2, we infer that low-level communities in the encoding tree provide a more faithful representation of the graph’s topology, as they simultaneously preserve local connectivity and global structural constraints. Accordingly, for each leaf node $\alpha$, we assign its associated low-level community $\mathcal{C}^{\ell}_{\alpha^{+}}$ as a topological context to enhance the accuracy of downstream inference by the LLMs. The proof process of Theorems 1 and  2 is provided in Appendix F.1-F.2.
@@ 4.2 Tree-prompted LLM Inference
Reception-aware Leaf Augmentation. To construct effective in-context prompts for LLMs, the quality of node-specific text is essential. Thus, we perform tree-guided text propagation and aggregation within each low-level community $\mathcal{C}^{\ell}$ to achieve leaf augmentation for $\epsilon$-guaranteed $t^{\star}$. The key idea is that nodes in low-level communities typically show high connectivity and homophily, enabling efficient text sharing. Using this property, we filter and combine relevant texts for each leaf node based on a similarity threshold $\epsilon$, helping to reduce noise. This process is formally defined as:
$t^{\star}_{\alpha}=\operatorname{Concat}\left(\left\{t_{\alpha}\}\cup\{t_{j}\mid j\in\text{Top}_{k}\left(\left\{w_{\alpha\beta}\mid\beta\in\mathcal{C}^{\ell}_{\alpha^{+}}\right\}\right),w_{\alpha\beta}\geq\epsilon\right\}\right),$ (2)
$w_{\alpha\beta}=\frac{\exp\left(\operatorname{sim}(x_{\alpha},x_{\beta})\right)}{\sum\limits_{\gamma\in\mathcal{C}^{\ell}_{\alpha^{+}}}\exp\left(\operatorname{sim}(x_{\alpha},x_{\gamma})\right)},\>\>\>\>\operatorname{sim}(x_{\alpha},x_{\beta})=\frac{\sum_{i=1}^{f\text{-dim}}x_{\alpha}^{i}\cdot x_{\beta}^{i}}{\sqrt{\sum_{i=1}^{f\text{-dim}}(x_{\alpha}^{i})^{2}\sum_{i=1}^{f\text{-dim}}(x_{\beta}^{i})^{2}}}.$ (3)
Community of Thought. Building on the enhanced leaf nodes, we propose a Community of Thought (CoT) prompt mechanism, where each node is represented together with its community-aware descriptions and related neighbors. This design enables LLMs to jointly capture topological and semantic information, thereby supporting more reliable inference. Through this mechanism, the LLM can (i) infer the probability distribution over label classes for each node, and (ii) generate high-quality embeddings within its semantic space. These embeddings are then employed for fine-grained optimization of the structural encoding tree. In summary, CoT facilitates efficient in-context learning for LLMs by leveraging the structure-aware encoding tree. Formally, given a target leaf node $\alpha$, the LLM inference with CoT can be expressed as:
$\mathbf{z}_{\alpha}=\Psi\!\left(\mathrm{LLM\mbox{-}CoT}\!\left(\mathcal{D}_{\text{task}},\mathcal{D}_{\text{topo}},\mathcal{D}_{\text{sema}},t^{\star}_{\alpha}\right)\right),\quad\Psi:\text{ANS}_{\alpha}\mapsto\mathbb{R}^{N_{\text{class}}},$ (4)
$y^{cls}_{\alpha}\sim\mathrm{Categorical}\!\left(p_{\theta}(y\mid\mathbf{z}_{\alpha})\right),\quad p_{\theta}(y=i\mid\mathbf{z}_{\alpha})=\frac{\exp(z_{\alpha,i})}{\sum_{j=1}^{N_{\text{class}}}\exp(z_{\alpha,j})},\;\;i=1,\ldots,N_{\text{class}},$ (5)
where $\Phi(\cdot)$ is the extraction function that maps the LLM’s textual answer $\text{ANS}_{\alpha}$ into a numerical logit vector $\mathbf{z}_{\alpha}\in\mathbb{R}^{N_{\text{class}}}$, $y^{cls}_{\alpha}\in\Delta^{N_{\text{class}}-1}$ denotes the soft label probability simplex for node $\alpha$. Here, $\mathcal{D}_{\text{task}}$ specifies the task, while $\mathcal{D}_{\text{topo}}$ and $\mathcal{D}_{\text{sema}}$ encode topological and semantic context, respectively. An illustration of prompt construction is provided in Fig. 2.
@@
Leaf Dependency Allocation. While the structural encoding tree captures graph topology, effective GSL also requires incorporating text-driven node attributes. To this end, we optimize low-level communities by reallocating leaf dependencies under the guidance of LLM-derived soft labels. Specifically, we (i) initialize clusters from the original structural encoding tree, (ii) reassign minority-label nodes to align with the dominant class, and (iii) adaptively refine cluster granularity using the silhouette coefficient to form communities with stronger homophily. This yields the optimized structural encoding tree $\mathcal{T}^{\star}$, which serves as the foundation for tree sampling (see Fig. 2). The procedure can be compactly expressed as:
$\mathcal{P}^{\star}=\arg\max_{\mathcal{P}}\Bigg\{\sum_{\mathcal{C}\in\mathcal{P}}\max_{c}\sum_{v\in\mathcal{C}}y^{cls}_{v,c}+\beta\,\sum_{\mathcal{C}\in\mathcal{P}}\mathrm{Sil}(\mathcal{C};\mathbf{Y}^{cls})\Bigg\},\;\;\mathcal{T}^{\star}=\mathrm{UpdateTree}(\mathcal{T},\mathcal{P}^{\star}).$ (6)
Here, $\mathcal{P}$ denotes a partition of the leaf nodes into communities, $y^{cls}_{v,c}$ is the probability of node $v$ belonging to class $c$, and $\mathrm{Sil}(\mathcal{C};\mathbf{Y}^{cls})$ is the silhouette coefficient evaluating the homophily of community $\mathcal{C}$. The optimized partition $\mathcal{P}^{\star}$ is then used to update the structural encoding tree $\mathcal{T}$, yielding the refined tree $\mathcal{T}^{\star}$. The adaptive clustering algorithm is detailed in Appendix E.2.